# How to Leverage Digit Embeddings to Represent Numbers?

### v1

@@ 4 Integrating Aggregated Embeddings
In contrast, the second approach aims to refine the model’s output by improving how numbers are represented in the learned outcomes. This is achieved by incorporating the aggregation in the loss function, encouraging the model to generate number embeddings that align more closely to the correct numerical values. Specifically, this method includes an additional term in the loss calculation, which accounts for the distance between the aggregated embedding of the predicted numbers and that of the true numbers. This targeted intervention is particularly effective in tasks requiring precise numerical predictions, helping the model develop a more nuanced and accurate representation of numbers.
@@ 4.2 Aggregation in Loss Function
Language generation models typically use a cross-entropy loss function ($\mathcal{L}_{CE}$) (Lewis et al., 2020; Raffel et al., 2020). To improve the model’s ability to predict numbers accurately, we introduce an auxiliary loss ($\mathcal{L}_{AUX}$) to calculate the mean squared error between the aggregate embedding of the gold and predicted numbers. Understanding and predicting numbers is inherently more complex than predicting a single word or sub-word because they consist of multiple digits, each carrying different significance. For example, in answering the question “Mary’s salary is £900 a month, but she pays £579 in rent. How much salary does she have left at the end of each month?”, the answers 320, 230, 32, or 456 are all incorrect. However, 320 is more accurate compared to others because its magnitude is closer to the correct answer, 321. Incorporating this new auxiliary loss would help the model predict digits that are closer to the gold answer, enhancing its precision in numerical predictions by recognising the relative significance of each digit within a number.
@@
The two losses are linearly interpolated by a hyperparameter, $\lambda$:
@@
Table 1: Results change from baseline after including aggregate embeddings in input embedding ([AGG] + Digits) and auxiliary loss (Digits + Aux Loss) for BART base, FLAN base and FLAN large. Darker shades of green and red indicate an absolute change greater than 1%.
@@ 5 Experimental Setup
Both methods are evaluated on two different pretrained models, BART base (140M) Lewis et al. (2020) and FLAN base (250M) Wei et al. (2022a). Additionally, we evaluate on FLAN large (780M) to explore the effect of model size. All of these models are encoder-decoders. BART is pre-trained on five corrupted document tasks from books and Wikipedia data. FLAN is an instruction-finetuned version of T5 Raffel et al. (2020) which is trained on C4 using transfer learning.
We evaluate our proposed methods on two different test sets: FERMAT Sivakumar and Moosavi (2023), and MAWPS Koncel-Kedziorski et al. (2016). Both FERMAT and MAWPS consist of English maths worded problem that can be tackled by BART and FLAN as shown by Sivakumar and Moosavi (2023) and where the answer is a single number. This enables us to evaluate our method strictly on numerical outputs reducing the interference of other difficulties such as predicting words and units, or extracting spans. FERMAT is a multi-view evaluation set which has different test sets with different number representations while keeping the maths problem fixed. The different test sets distinguish different number types of which we select the ones that separate integers into number lengths, mix integers less than 1000, mix integers greater than 1000, one and two decimal place numbers, and a test set scaled up to more than 4-digit numbers; these allow us to evaluate which number representation the models support better. FERMAT’s training set is augmented from templates making it independent to its test sets. MAWPS, on the other hand, has the same domain for both training and testing. It is a widely used dataset to evaluate numerical reasoning, chiefly because it is small and easy to train with small models. We finetune the models on each dataset’s respective training data (see Appendix B) using the hyperparameters described in Appendix C.
@@ 6 Impact of Integrating Aggregations
FLAN large, on the other hand, has a more balanced performance but an overall higher improvement when the aggregation is incorporate in the input as shown particularly from all the green cells in the row [AGG] + Digits. Therefore, a certain model size may be required to learn a new token and leverage the information it provides. This reinforces that an aggregated embedding provides useful signal to improve number understanding but how it is integrated is also crucial.
@@
For the 1000+ columns, using accuracy, the pattern is not evident, however, from Appendix D, using the auxiliary loss clearly reduces the CER more than explicitly using the aggregation in the input. The auxiliary loss encourages the model to predict the correct answer as the CER is lower. However, since the weights assigned to each digit position is lower as it gets closer to the units, the auxiliary accounts less for it, reducing precision. As a consequence, despite the CER reducing, since the entire number is not predicted correctly, improvement fails to be reflect in the accuracy.
7 Analysis of Aggregation Embedding in the Input
@@
Additionally, we test the impact of providing the aggregated token by replacing it with a randomly initialised [PAUSE] token akin to Goyal et al. (2024). From Table 2, we observe that for BART, nor [AGG], nor [PAUSE] have a great positive impact on the performance. This confirms that BART struggles to learn new tokens from finetuning alone. The FLAN models are more adaptable to the new tokens as seen by the greener rows. However, the overwhelming bold entries with the [PAUSE] token indicate that both FLAN base and large perform better with a [PAUSE] token acting as a blank space for the model to process the information. It may also be that the model uses this token to create an implicit representation of the number. Nevertheless, the average improvement between the [PAUSE] and [AGG] differs by less than 0.5% implying that a different aggregation function or a full hyperparameter search could reverse the trend.
@@ 8 Future Work
Our proposed aggregation strategy has shown encouraging steps towards better number representation. However, as with observation made in previous work, the effect of new strategies report minimal improvement on smaller models but greater impact on larger models Cobbe et al. (2021); Wei et al. (2022b). Therefore, an evaluation of our proposed method on larger scale models would verify the scalability of this approach.
The weighting scheme, presented in Equation 1, offers a straightforward method for aggregating digit embeddings. However, as numbers increase in length, their aggregated embeddings tend to drift away from the original numerical embedding space. This divergence could be addressed by enabling the model to adapt to this new embedding space by exploring extended pretraining, or alternative weighting schemes that remain closer to the numerical subspace while satisfying the criteria outlined in Section 3.
@@ 9 Conclusion
Improving numerical reasoning is a challenging task, increasing model sizes or focusing on data augmentation helps but at the cost of a substantial additional training time or computations. Digit tokenisation has been a pioneering work in improving how models encode and decode numbers, however the aggregation of the digit is done implicitly. We advance this idea by explicitly providing an aggregated number embedding that is more mathematically sound. These embeddings are generated as weighted sums of the digit embeddings by accounting for the digits relative position in the number. We then incorporate them in two model agnostic forms: in the input level as an additional token, and in an auxiliary MSE loss. Our promising results demonstrate that, as a proof-of-concept, even a straightforward aggregation with simple incorporation techniques can positively impact number understanding. Therefore, testing it at larger scale, developing sophisticated aggregation functions, and refining the integration of the auxiliary loss presents valuable avenues for future research.
@@ 10 Limitations
Some of the limitations of this work is discussed in the Future Work section. However, we give detail of more limitations relating to the size of the models used, and the compatibility and growth of our proposed weighted aggregation function.
@@
In addition, while the weighted scheme is designed using mathematical priors, it is specifically created for integers, therefore it may not be compatible with decimals or alternative representation of numbers such as 01 for 1. Nonetheless, from Table 5, we note that CER reduces for both 1dp and 2dp therefore our aggregated embedding method has promising scope for all numbers. Lastly, the weights function described in Equation 1 does not converge, therefore for a sufficiently large number of digit it would grow beyond the accuracy provided by the model. However, we explain in Section 3 with the aid of Figure 2 that, for up to 6-digits, the weighted scheme functions well with no signs of deterioration. Moreover, in natural text, very large numbers tend to be shorten using a more appropriate unit, for example, the world population of 8114693010 is more often expressed as 8 billion reducing the numbers of digits needed considerably.
@@ Acknowledgements
This work was supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation [grant number EP/S023062/1]. Additional thanks to Danae Sanchez Villegas for her continued feedback in this research.
@@ Appendix C Hyperparameters
Appendix D Character Error Rate (CER) Results
FERMAT
Incorporating Weights (CER %) MAWPS
Original
Commuted
Integers 0 to 1000
2-digit integers
3-digit integers
4-digit integers
1000+
1000+ same
1dp random
2dp random
a+b
a-b
a*b
a/b
Table 5 presents the character error rate (CER) for incorporating the weighted aggregation as an input token and in the auxiliary loss, for all three models.

### v2

@@ 4 Integrating Aggregated Embeddings
In contrast, the second approach aims to refine the model’s output by improving how numbers are represented in the learned outcomes. This is achieved by incorporating the aggregation in the loss function, encouraging the model to generate number embeddings that align more closely to the correct numerical values. Specifically, this method includes an additional term in the loss calculation, which accounts for the distance between the aggregated embedding of the predicted number and that of the true number. This targeted intervention is particularly effective in tasks requiring precise numerical predictions, helping the model develop a more nuanced and accurate representation of numbers.
@@ 4.2 Aggregation in Loss Function
Language generation models typically use a cross-entropy loss function ($\mathcal{L}_{CE}$) (Lewis et al., 2020; Raffel et al., 2020). To improve the model’s ability to predict numbers accurately, we introduce an auxiliary loss ($\mathcal{L}_{AUX}$) to calculate the mean squared error between the aggregate embedding of the gold and predicted numbers. Understanding and predicting numbers is inherently more complex than predicting a single word or sub-word because they consist of multiple digits, each carrying different significance. For example, in answering the question “Mary’s salary is £900 a month, but she pays £579 in rent. How much salary does she have left?”, the answers 320, 230, 32, or 456 are all incorrect. However, 320 is more accurate compared to others because its magnitude is closer to the correct answer, 321. Incorporating this new auxiliary loss would help the model predict digits that are closer to the gold answer, enhancing its precision in numerical predictions by recognising the relative significance of each digit within a number.
@@
Finally, the two losses are linearly interpolated by a hyperparameter, $\lambda$:
@@
Table 1: Results change in Accuracy from baseline after including aggregate embeddings in input embedding ([AGG] + Digits) and auxiliary loss (Digits + Aux Loss) for BART base, FLAN base and FLAN large. Darker shades of green and red indicate an absolute change greater than 1%.
FERMAT
Incorporating Weights (CER %) MAWPS
Original
Commuted
Integers 0 to 1000
2-digit integers
3-digit integers
4-digit integers
1000+
1000+ same
1dp random
2dp random
a+b
a-b
a*b
a/b
@@ 5 Experimental Setup
Both methods are evaluated on two different pretrained models, BART base (140M) Lewis et al. (2020) and FLAN base (250M) Wei et al. (2022a). Additionally, we evaluate on FLAN large (780M) to explore the effect of model size. BART is an encoder-decoder pre-trained on five corrupted document tasks from books and Wikipedia data. FLAN is an instruction-finetuned version of T5 Raffel et al. (2020) which is trained on C4 using transfer learning.
We evaluate our proposed methods on two different test sets: FERMAT Sivakumar and Moosavi (2023), and MAWPS Koncel-Kedziorski et al. (2016). Both FERMAT and MAWPS consist of English maths worded problem that can be tackled by BART and FLAN as shown by Sivakumar and Moosavi (2023) and where the answer is a single number. This enables us to evaluate our method strictly on numerical outputs reducing the interference of other difficulties such as predicting words and units, or extracting spans. FERMAT is a multi-view evaluation set which has different test sets with different number representations while keeping the maths problem fixed. The different test sets distinguish different number types of which we select the ones that separate integers into digit length (2-digit, 3-digit, 4-digit), contain a mixture of integers less than 1000, contain a mixture of integers greater than 1000, the sets of one and two decimal place numbers, and a test set that takes the original set and scales the number to more than 4-digit numbers; these allow us to evaluate which number representation the models support better. FERMAT’s training set is augmented from different templates making it independent to its test sets. MAWPS, on the other hand, has the same domain for both training and testing. It is a widely used dataset to evaluate numerical reasoning, chiefly because it is small and easy to train with small models. We finetune the models on each dataset’s respective training data (see Appendix B) using the hyperparameters described in Appendix C.
@@ 6 Impact of Integrating Aggregations
FLAN large, on the other hand, has a more balanced performance but an overall higher improvement when the aggregation is incorporated in the input as shown particularly from all the green cells in the row [AGG] + Digits. Therefore, a certain model size may be required to learn a new token and leverage the information it provides. This reinforces that an aggregated embedding provides useful signal to improve number understanding but how it is integrated is also crucial.
For the operations, the improvements is generally positive across all of them, however, evidently greater for addition and subtraction than multiplication and division. This resonates with the fact that digits positions are more informative for the first two operations, especially when, for instance, aligning them to perform calculations.
@@
For the 1000+ columns, using accuracy, the pattern is not evident, however, Table 2 presents the character error rate (CER) comparing both incorporating strategies for all three models, and highlights that using the auxiliary loss clearly reduces the CER more than explicitly using the aggregation in the input. The auxiliary loss encourages the model to predict the correct answer as the CER is lower. However, since the weights assigned to each digit position is lower as it gets closer to the units, the auxiliary accounts less for it, reducing precision. As a consequence, despite the CER reducing, since the entire number is not predicted correctly, improvement fails to be reflected in the accuracy.
7 Analysis of Aggregation Embedding in the Input
@@
Additionally, we test the impact of providing the aggregated token by replacing it with a randomly initialised [PAUSE] token akin to Goyal et al. (2024). From Table 3, we observe that for BART, neither [AGG], nor [PAUSE] have a great positive impact on the performance. This confirms that BART struggles to learn new tokens from finetuning alone. The FLAN models are more adaptable to the new tokens as seen by the greener rows. However, the overwhelming bold entries with the [PAUSE] token indicate that both FLAN base and large perform better with a [PAUSE] token acting as a blank space for the model to process the information. It is possible that that the model uses this token to create an implicit representation of the number. Nevertheless, the average improvement between the [PAUSE] and [AGG] differs by less than 0.5% implying that a different aggregation function or a full hyperparameter search could reverse the trend.
@@ 8 Future Work
Our proposed aggregation strategy has shown encouraging steps towards better number representation. However, as with observations made in previous work, the effect of new strategies report minimal improvement on smaller models but greater impact on larger models Cobbe et al. (2021); Wei et al. (2022b). Therefore, an evaluation of our proposed method on larger scale models would verify the scalability of this approach.
The weighting scheme, presented in Equation 1, offers a straightforward method for aggregating digit embeddings. However, as numbers increase in length, their aggregated embeddings tend to drift away from the original numerical embedding subspace. This divergence could be addressed by enabling the model to adapt to this new embedding space by exploring extended pretraining, or constructing weighting schemes that remain closer to the numerical subspace while satisfying the criteria outlined in Section 3.
@@ 9 Conclusion
Improving numerical reasoning is a challenging task, increasing model sizes or focusing on data augmentation helps but at the cost of a substantial additional training time or computations. Digit tokenisation has been a pioneering work in improving how models encode and decode numbers, however the aggregation of the digit is done implicitly. We advance this idea by explicitly providing an aggregated number embedding that is more mathematically sound. These embeddings are generated as weighted sums of the digit embeddings by accounting for the digits relative position in the number. We then incorporate them in two model agnostic forms: in the input level as an additional token, and in an auxiliary MSE loss. Our promising results demonstrate that, as a proof-of-concept, even a straightforward aggregation with simple incorporation techniques can positively impact number understanding. Therefore, testing it at a larger scale, developing sophisticated aggregation functions, and refining the integration of the auxiliary loss presents valuable avenues for future research.
@@ 10 Limitations
Some of the limitations of this work are discussed in the Future Work section. However, we give details of further limitations relating to the size of the models used, and the compatibility and growth of our proposed weighted aggregation function.
@@
In addition, while the weighted scheme is designed using mathematical priors, it is specifically created for integers, therefore it may not be compatible with decimals or alternative representation of numbers such as 01 for 1. Nonetheless, from Table 2, we note that CER reduces for both 1dp and 2dp therefore our aggregated embedding method has promising scope for all numbers.
Furthermore, the weight function described in Equation 1 does not converge, therefore for a sufficiently large number of digit it would grow beyond the accuracy provided by the model. However, we explain in Section 3 with the aid of Figure 2 that, for up to 6-digits, the weighted scheme functions well with no signs of deterioration. Moreover, in natural text, very large numbers tend to be shorten using a more appropriate unit, for example, the world population of 8114693010 is more often expressed as 8 billion reducing the numbers of digits needed considerably. But this raises the question of predicting the correct unit which would lead to future work.
Nonetheless, our weighting scheme leverages digit embedding, therefore it is heavily dependent on them, particularly on the relative distance of the digit embedding to one another. In FLAN large, the embedding of the digit 0 is more distant from the embeddings of other digits, which causes it to frequently include numbers ending in 0 when the target contains a 0, or to exclude them otherwise. As explained in Section 3, the optimal neighbour would include all numbers with different unit value, and this alone would achieve an F1-score of at least 70%.
Lastly, the experiments were conducted using a single random seed. While this ensures consistency and reproducibility, having access to better resources would have enabled us to run the experiments with multiple seeds. This would have allowed us to calculate the average improvement achieved by using aggregated digit embeddings to represent numbers.
@@ Acknowledgements
This work was supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation [grant number EP/S023062/1]. We also acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing. Additional thanks to the reviewers for their encouraging comments and discussion, and particularly to Danae Sanchez Villegas, Mugdha Pandya, Valeria Pastorino, Huiyin Xue and Constantinos Karouzos for their continued feedback in through the research.